A newspaper article estimated that the mean cost for preventive dental care was $165 per year. They indicated that the estimate was based on a sample of 100 people and that the margin of error was 2.8. Assuming a 95% confidence level was used, calculate the value of the standard error.


1
Expert's answer
2021-10-03T17:20:42-0400

Given "\alpha=0.05, df=n-1=100-1=99" degrees of freedom.

The critical t-value is "t_c=1.984217."

The margin of error is "ME=t_c\times\dfrac{s}{\sqrt{n}}=2.8."

Then the standard error is


"SE=\dfrac{s}{\sqrt{n}}=\dfrac{ME}{t_c}=\dfrac{2.8}{1.984217}=1.411"